OpenStudy (anonymous):

Assume that the car lot contains 20% Volvos, 30% Porches, and 50% Lincolns. Of the Volvos, 60% have 2 airbags, 40% of the Porches have 2 airbags, and 20% of the Lincolns have airbags. Furthermore, 60% of the Volvos, 40 % of the Porches, and 90% of the Lincolns are white. The property of being white is independent from having 2 airbags. You are assigned a car at random. If the car has 2 airbags and is white, what is the probability that it is a Volvo_____________?

OpenStudy (dumbcow):

30.69 %

OpenStudy (anonymous):

that translates to .3069 right?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

that wasnt right

OpenStudy (dumbcow):

ok i'll double check...is there any rounding required when you put in an answer?

OpenStudy (anonymous):

no

OpenStudy (dumbcow):

i don't see any mistakes on my end P(airbag) = 0.34 P(white) = 0.69 P(both) = 0.2346 P(volvo and both) = 0.36*.2 = 0.072 P(V| both) = .072/.2346 = 0.3069054

OpenStudy (anonymous):

still wont work

OpenStudy (dumbcow):

oh well...hopefully someone will correct me :)

OpenStudy (anonymous):

thank you anyway this stuff has killed me you helped me out greatly

OpenStudy (dumbcow):

np

OpenStudy (dumbcow):

update: i noticed my mistake A= 2 airbags, W=white V = volvo, X = porche, L = lincoln P(AW) = P(AW|V)*P(V) + P(AW|X)*P(X) + P(AW|L)*P(L) = (.36*.2) + (.16*.3) +(.18*.5) = 0.21 P(VAW) = P(AW|V)*P(V) = .36 * .2 = 0.072 P(V |AW) = P(VAW)/P(AW) = .072/.21 = 0.34286